A tin containing 5000cm>3 of paint has a mass of 7kg. If the mass of the rmpty tin including the lid is 0.5kg calculate the density of the paint in g/cm>3.
If the tin is made of metal which has a density of 7800kg/m>3 calculate the volume used to make tin and lid.
1
Expert's answer
2019-10-22T10:24:32-0400
"\rho _p=\frac{M-m}{V}=\frac{7000-500}{5000}=1.3\frac{g}{cm^3}"

The volume used to make tin and lid:


"V=\frac{m}{\rho_m}=\frac{0.5}{7800}=6.4\cdot 10^{-5}\ m^3"